Derive the eqn : dG=VdP-sdT..................................

Grade:11

2 Answers

Vikas TU
7 years ago
Diffrentiate:  G = H - TS
we get, dG = dH - d(TS) = dH - TdS - SdT. 
After, differentiating H = U + pV
we get, dH = dU + pdV + Vdp.
Also, dU = dq + dw = dq - pdV,
so dH = dq + Vdp.
hence, we get the final eqn. as
: dG = dq + Vdp - TdS - SdT
 
S S
7 years ago
The first answer is correct throughout except for the last part....We know, that dS=dq/THence the final equation becomesdG=VdP-SdT+dq-Tdq/T=>dG=VdP-SdT .